# On a fast consistent selection of nested models with possibly unnormalised probability densities

### v2

@@ 5.2 Car Data
The data plot suggests a nonlinear relationship between mpg and horsepower. Gareth et al. (2013) fitted a polynomial regression model with Gaussian errors to this dataset. In this study, we apply a polynomial regression model with $N\times t$ errors and use the T-GIC1 and T-GIC2 to select the degree of the polynomial, setting candidate degrees from 1 to 10. For comparison, we also use AIC and BIC with Gaussian errors. For $N\times t$ errors we employ the model in equation (4.7), and for Gaussian errors we employ the model given by
@@
Tables 13 and 14 report model selection and parameter estimation, respectively. Under Gaussian errors, AIC selects a seventh-degree polynomial regression model, while BIC selects a quadratic polynomial regression model. In contrast, both T-GIC1 and T-GIC2 select a quadratic polynomial regression model with $N\times t$ errors, and give estimates similar to those based on the BIC with Gaussian errors. This suggests that a quadratic model chosen by BIC is an appropriate model rather than the seventh-degree model chosen by AIC. This example shows that T-GIC1 and T-GIC2 can be used to check the choice by AIC and BIC, even when the latter can be applied, and highlights an additional utility of T-GIC coupled with the $N\times t$ PDF.
@@
$p=2$ (T-GIC1, T-GIC2) (-0.3838, 0.0800) 3.0288 0.3757 0.4973 3.2389
We also examine the skewness and excess kurtosis of the residuals and consider the matching between the theoretical values and their sample estimates based on the fitted residuals, as shown in Table 15. Approximate standard errors (SE) of the sample estimates are obtained by a Bootstrap procedure, with 1000 replications. For the fitted models with Gaussian errors, the skewness and the excess kurtosis of the errors are both theoretically 0. On the other hand, for $N\times t(\alpha,k)$ errors, the skewness is theoretically 0, but the excess kurtosis would not be 0 unless $k=0$. Since it is difficult to calculate the excess kurtosis for $N\times t(\alpha,k)$ PDF due to its intractable normalizing constant, we sample from $N\times t$ PDF and use the mean of the sample estimates from 1000 replications as the theoretical value, along with its SE. Apparently, the fitted residuals of the models chosen by AIC, BIC, T-GIC1 and T-GIC2 have all produced very small negative skewness of similar size, matching their theoretical value of zero reasonably well. However, for excess kurtosis, while the matching is far from being satisfactory for the Gaussian models chosen by AIC and BIC, the matching for models chosen by T-GIC1 and T-GIC2 coupled with $N\times t$ PDFs is much better.
@@
Noise Model Skewness Excess kurtosis
@@
$N(0,\sigma^{2})$ AR(7) 0 -0.08 (0.22) 0 1.13 (0.59)
AR(2) 0 -0.14 (0.20) 0 0.87 (0.49)
$N\times t(\alpha,k)$ AR(2) 0 -0.22 (0.23) 1.29 (0.21) 1.33 (0.54)
@@ 5.3 Wind Direction Data
Finally, we fit a bivariate model with von Mises PDF, as described in section 4.3, to some wind direction data, using T-GIC1 and T-GIC2 for model selection. Here, the wind direction is represented as a circular variable in radians. Matsuda et al. (2021) applied this model to wind direction data from Tokyo at 00:00 and 12:00 in 2008. For the sake of cross-validation, we analyze more recent wind direction data from Tokyo at 00:00 ($x_{1}$) and 12:00 ($x_{2}$) over 365 days in 2023, obtained from the Japan Meteorological Agency website. The data are discretized into 16 bins, such as north-northeast. Figure 3 presents the corresponding 2-d histogram.
@@
We fit the data using two candidate models, $m_{1}$ and $m_{2}$, representing two scenarios: $\lambda=0$ and $\lambda\in R$ in equation (4.8), respectively. For comparison, results for T-GIC1, T-GIC2 are summarized in Table 16. For $m_{2}$, both T-GIC1 and T-GIC2 are higher, indicating a better fit than $m_{1}$. This suggests that the wind directions at Tokyo on 00:00 and 12:00 in 2023 are dependent, consistent with the results in 2008 by Matsuda et al. (2021).
Table 16: Parameter estimate and T-GIC results for models $m_{1}$ and $m_{2}$.
Model Order $\kappa_{1}$ $\kappa_{2}$ $\mu_{1}$ $\mu_{2}$ $\lambda$ T-GIC1 T-GIC2
@@ 6 Conclusion
In this paper, based on GIC (Cheng and Tong, 2024), we propose a consistent selection criterion, T-GIC, for nested models with possibly unnormalized PDFs. Compared with existing model selection methods, T-GIC offers three advantages. First, it can handle both normalized and unnormalized PDFs. Second, it enjoys selection consistency for the case of a finite sequence of strictly nested models under mild regularity conditions. Third, it reduces significantly computational costs by avoiding the calculation of the normalizing constant and the bias correction. To showcase the efficacy of T-GIC for unnormalized PDFs, we have studied AR models and polynomial regression models with $N\times t$ noise/errors, the normalizing constants of which are typically intractable. Extensive simulation studies and real data applications have demonstrated consistency and effectiveness of T-GIC. Furthermore, we have shown excellent performance of T-GIC on PDFs with bounded support through experiments on models with bivariate von Mises PDFs and real wind direction data.
We have discussed how to use T-GIC for consistent model selection with PDFs supported on $\mathbb{R}^{d}.$ The simulation results and the real example with wind data strongly suggest that a promising direction for future research is to extend T-GIC to cover PDFs supported on bounded intervals $(a,b)$. Another direction of extension is to discrete data, thereby availing T-GIC of the opportunity of selecting an appropriate model, such as an Ising model (Friel, 2013; Everitt et al., 2017), in the area of discrete Markov random fields and spatial statistics. It is also intriguing to explore the use of combinations of different multiplying factors $C(n,k)$. Drawing inspiration from AIC and BIC, we have set $C(n,k)$ to $\exp\{-2\#(M_{k})/n\}$ and $n^{-\#(M_{k})/n}$, respectively. Although $\exp\{-2\#(M_{k})/n\}$ may not yield a consistent estimate of the true order, there is significant scope for combining it with $n^{-\#(M_{k})/n}$, similar to existing approaches for combining AIC and BIC in model selection (Hsu et al., 2019). Last, but not least, an exploration of T-GIC into non-nested models should be exciting.
Appendix A Appendix
A.1 Proof of Proposition 1
@@
00 0 $=P_{n}\nabla_{\alpha}W(x,p_{M(\hat{\alpha}_{0},\beta^{*})})$
$=P_{n}\nabla_{\alpha}W(x,p_{M(\alpha^{*},\beta^{*})})+P_{n}\nabla_{\alpha}\nabla^{T}_{\alpha}W(x,p_{M(\alpha^{*},\beta^{*})})(\hat{\alpha}_{0}-\alpha^{*})+O_{p}(n^{-1})$ (A.1)
@@
$+P_{n}\nabla_{\alpha}\nabla^{T}_{\beta}W(x,p_{M(\alpha^{*},\beta^{*})})(\hat{\beta}-\beta^{*})+O_{p}(n^{-1}).$ (A.2)
@@
$\hat{\alpha}_{0}-\alpha^{*}=\hat{\alpha}-\alpha^{*}+D^{-1}_{\alpha^{*},\alpha^{*}}D_{\alpha^{*},\beta^{*}}(\hat{\beta}-\beta^{*})+O_{p}(n^{-1}),$ (A.3)
@@
$=\begin{pmatrix}(\hat{\alpha}-\alpha^{*})^{T}&(\hat{\beta}-\beta^{*})^{T}\end{pmatrix}\begin{pmatrix}D_{\alpha^{*},\alpha^{*}}&D_{\alpha^{*},\beta^{*}}\\ D_{\beta^{*},\alpha^{*}}&D_{\beta^{*},\alpha^{*}}D^{-1}_{\alpha^{*},\alpha^{*}}D_{\alpha^{*},\beta^{*}}\end{pmatrix}\binom{\hat{\alpha}-\alpha^{*}}{\hat{\beta}-\beta^{*}}+O_{p}(n^{-3/2})$ (A.4)
@@
$=\begin{pmatrix}(\hat{\alpha}-\alpha^{*})^{T}&(\hat{\beta}-\beta^{*})^{T}\end{pmatrix}\begin{pmatrix}D_{\alpha^{*},\alpha^{*}}&D_{\alpha^{*},\beta^{*}}\\ D_{\beta^{*},\alpha^{*}}&D_{\beta^{*},\beta^{*}}\end{pmatrix}\binom{\hat{\alpha}-\alpha^{*}}{\hat{\beta}-\beta^{*}}+O_{p}(n^{-3/2})$ (A.5)
@@
$D(\theta^{*})=\begin{pmatrix}D(\alpha^{*},\alpha^{*})&D(\alpha^{*},\beta^{*})\\ D(\beta^{*},\alpha^{*})&D(\beta^{*},\beta^{*})\end{pmatrix}$ (A.6)
Subtracting (A.5) from (A.4) yields:
@@
$=-(\hat{\beta}-\beta^{*})^{T}(D_{\beta^{*},\beta^{*}}-D_{\beta^{*},\alpha^{*}}D^{-1}_{\alpha^{*},\alpha^{*}}D_{\alpha^{*},\beta^{*}})(\hat{\beta}-\beta^{*})+O_{p}(n^{-3/2})$ (A.7)
However, $\log P_{n}\{W(x,p_{M(\hat{\alpha},\hat{\beta})})\}-\log P_{n}\{W(x,p_{M(\hat{\alpha}_{0},\beta^{*})})\}=\log GIC(k)-\log GIC(k_{0}).$ Recall $\sqrt{n}(\hat{\theta}-\theta^{*})$ is asymptotically normally distributed with mean zero and covariance matrix equal to $D^{-1}(\theta^{*})\Lambda(\theta^{*})D^{-T}(\theta^{*})$. It follows from routine algebra that $\sqrt{n}(\hat{\beta}-\beta^{*})$ is asymptotically normal with zero mean vector and covariance matrix equal to $B(\theta^{*})B^{T}(\theta^{*})$. Hence, $n\times[\log GIC(k)-\log GIC(k_{0})]$ converges in distribution to $Z^{T}A(\theta^{*})Z$.
A.2 Proof of Theorem 1
@@
First, we observe that for any model $M$ satisfying Assumptions 6-7 and assuming the validity of the law of large numbers, then it follows from routine analysis that $GIC(M(\theta))=P_{n}\{W(x,p_{M(\theta)})\}$ converges uniformly in probability to its population version $GIC_{\infty}(M(\theta))=P\{W(x,p_{M(\theta)})\}$. It follow from Assumption 7 that $GIC_{\infty}(M_{k}(\theta_{k}))$ is a Lipschitz-continuous, as a function of the parameter $\theta_{k}$, hence it attains its maximum value, denoted by $\mathcal{M}_{k}$, owing to the compact parameter space assumption (aka Assumption 6). Suppose $k_{0}$ is the smallest $k$ such that $M_{k}$ contains the true model. Then, $\mathcal{M}_{k}<\mathcal{M}_{k_{0}}$ for all $1\leq k<k_{0}$, whereas $\mathcal{M}_{k}=\mathcal{M}_{k_{0}}$ otherwise. Note that from Proposition 3 in Cheng and Tong (2024), $\mathcal{M}_{k_{0}}=H_{G}(p^{*})>0$ where $p^{*}$ is the true population pdf. Therefore, if for any $k$, $C(n,k)\to 1$ as $n\to\infty$, then the maximum T-GIC model selection criterion will not select any $k<k_{0}$, in probability.
Henceforth, consider the case that $k\geq k_{0}$. Since $\mathcal{M}_{k_{0}}>0$, $GIC(k)$ is positive, in probability, i.e., $GIC(k)>0$ holds with probability approaching 1 as sample size increases without bound. For ease of exposition, we shall assume that $GIC(k)$ is positive. Let $k>k_{0}$ be fixed. Consider the increment $D=\log\{T\text{-}GIC(k)\}-\log\{T\text{-}GIC(k_{0})\}=\log C(n,k)-\log C(n,k_{0})+\log GIC(k)-\log GIC(k_{0})$. By Proposition 1, $n\times\{\log GIC(k)-\log GIC(k_{0})\}$ converges weakly to some non-negative distribution. Consequently, $D=\log C(n,k)-\log C(n,k_{0})+O_{p}(1/n)$ so that if $n\times\log\{C(n,k)/C(n,k_{0})\}\to-\infty$ as $n\to\infty$, $D$ is negative in probability for $k>k_{0}$. This completes the proof of the consistency of the proposed T-GIC model selection criterion.

### v3

@@ 5.2 Car Data
The data plot suggests a nonlinear relationship between mpg and horsepower. James et al. (2013) fitted a polynomial regression model with Gaussian errors to this dataset. In this study, we apply a polynomial regression model with Baker errors and use the MIC1 and MIC2 to select the degree of the polynomial, setting candidate degrees from 1 to 10. For comparison, we also use AIC and BIC with Gaussian errors. For Baker errors we employ the model in equation (4.7), and for Gaussian errors we employ the model given by
@@
Tables 12 and 13 report model selection and parameter estimation, respectively. Under Gaussian errors, AIC selects a seventh-degree polynomial regression model, while BIC selects a quadratic polynomial regression model. In contrast, both MIC1 and MIC2 select a quadratic polynomial regression model with Baker errors, and give estimates similar to those based on the BIC with Gaussian errors. This suggests that a quadratic model chosen by BIC is an appropriate model rather than the seventh-degree model chosen by AIC. This example shows that MIC1 and MIC2 can be used to check the choice by AIC and BIC, even when the latter can be applied, and highlights an additional utility of MIC coupled with the Baker PDF.
@@
$p=2$ (MIC1, MIC2) (-0.3838, 0.0800) 3.0288 0.3757 0.4973 3.2389
We also examine the skewness and excess kurtosis of the residuals and consider the matching between the theoretical values and their sample estimates based on the fitted residuals, as shown in Table 14. Approximate standard errors (SE) of the sample estimates are obtained by a Bootstrap procedure, with 1000 replications. For the fitted models with Gaussian errors, the skewness and the excess kurtosis of the errors are both theoretically 0. On the other hand, for Baker errors, the skewness is theoretically 0, but the excess kurtosis would not be 0 unless $k=0$. Since it is difficult to calculate the excess kurtosis for Baker PDF due to its intractable normalizing constant, we sample from Baker PDF and use the mean of the sample estimates from 1000 replications as the theoretical value, along with its SE. Apparently, the fitted residuals of the models chosen by AIC, BIC, MIC1 and MIC2 have all produced very small negative skewness of similar size, matching their theoretical value of zero reasonably well. However, for excess kurtosis, while the matching is far from being satisfactory for the Gaussian models chosen by AIC and BIC, the excess kurtosis from the fitted model chosen by MIC1 and MIC2 coupled with Baker PDF is 1.29 (SE: 0.21) versus its residual counterpart of 1.33 (SE: 0.54), which is much better.
@@
Noise Degree Skewness Excess kurtosis
@@
$N(0,\sigma^{2})$ $p=7$ 0 -0.08 (0.22) 0 1.13 (0.59)
$p=2$ 0 -0.14 (0.20) 0 0.87 (0.49)
$N\times t(\alpha,k)$ $p=2$ 0 -0.22 (0.23) 1.29 (0.21) 1.33 (0.54)
@@ 5.3 Wind Direction Data
Finally, we fit a bivariate model with von Mises PDF, as described in section 4.3, to some wind direction data, using MIC1 and MIC2 for model selection. Here, the wind direction is represented as a circular variable in radians. Matsuda et al. (2021) applied this model to wind direction data from Tokyo at 00:00 and 12:00 in 2008. For the sake of cross-validation, we analyze more recent wind direction data from Tokyo at 00:00 ($x_{1}$) and 12:00 ($x_{2}$) over 365 days in 2023, obtained from the Japan Meteorological Agency website. The data are discretized into 16 bins, such as north-northeast. Figure 5 presents the corresponding 2-d histogram.
@@
We fit the data using two candidate models, $m_{1}$ and $m_{2}$, representing two scenarios: $\lambda=0$ and $\lambda\in R$ in equation (4.8), respectively. For comparison, results for MIC1, MIC2 are summarized in Table 15. For $m_{2}$, both MIC1 and MIC2 are higher, indicating a better fit than $m_{1}$. This suggests that the wind directions at Tokyo on 00:00 and 12:00 in 2023 are dependent, consistent with the results in 2008 by Matsuda et al. (2021).
Table 15: Parameter estimate and MIC results for models $m_{1}$ and $m_{2}$.
Model Order $\kappa_{1}$ $\kappa_{2}$ $\mu_{1}$ $\mu_{2}$ $\lambda$ MIC1 MIC2
@@ 6 Conclusion
In this paper, we propose a consistent selection criterion, MIC, for nested models with possibly unnormalized PDFs. Compared with existing model selection methods, MIC offers three advantages. First, it can handle both normalized and unnormalized PDFs. Second, it enjoys selection consistency for the case of a finite sequence of strictly nested models under mild regularity conditions. Third, it reduces significantly computational costs by avoiding the calculation of the normalizing constant and the bias correction. To showcase the efficacy of MIC for unnormalized PDFs, we have studied AR models and polynomial regression models with Baker noise/errors, the normalizing constants of which are typically intractable. Extensive simulation studies and real data applications have demonstrated consistency and effectiveness of MIC. Furthermore, we have shown excellent performance of MIC on PDFs with bounded support through experiments on models with bivariate von Mises PDFs and real wind direction data.
We have discussed how to use MIC for consistent model selection with PDFs supported on $\mathbb{R}^{d}.$ The simulation results and the real example with wind data strongly suggest that a promising direction for future research is to extend MIC to cover PDFs supported on bounded intervals $(a,b)$. Another direction of extension is to discrete data, thereby availing MIC of the opportunity of selecting an appropriate model, such as an Ising model (Friel 2013, Everitt et al. 2017), in the area of discrete Markov random fields and spatial statistics. It is also intriguing to explore the use of combinations of different multiplying factors $C(n,k)$. We have set $C(n,k)$ to $\exp\{-2\#(M_{k})/n\}$ and $n^{-\#(M_{k})/n}$, respectively. Although $\exp\{-2\#(M_{k})/n\}$ may not yield a consistent estimate of the true order, there is significant scope for combining it with $n^{-\#(M_{k})/n}$, similar to existing approaches developed in Ing (2007). Last, but not least, an exploration of MIC into non-nested models should be exciting.
7 Disclosure Statement
We declare that we do not have any relevant financial or non-financial competing interests.
[…]
Proof.
@@ 7 Disclosure Statement
$0$ $=P_{n}\nabla_{\alpha}W(x,p_{M(\hat{\alpha}_{0},\beta^{*})})$
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
However, $\log P_{n}\{W(x,p_{M(\hat{\alpha},\hat{\beta})})\}-\log P_{n}\{W(x,p_{M(\hat{\alpha}_{0},\beta^{*})})\}=\log GIC(k)-\log GIC(k_{0}).$ Recall $\sqrt{n}(\hat{\theta}-\theta^{*})$ is asymptotically normally distributed with mean zero and covariance matrix equal to $D^{-1}(\theta^{*})\Lambda(\theta^{*})D^{-T}(\theta^{*})$. It follows from routine algebra that $\sqrt{n}(\hat{\beta}-\beta^{*})$ is asymptotically normal with zero mean vector and covariance matrix equal to $B(\theta^{*})B^{T}(\theta^{*})$. Hence, $n\times[\log GIC(k)-\log GIC(k_{0})]$ converges in distribution to $Z^{T}A(\theta^{*})Z$. ∎
[…]
@@
First, we observe that for any model $M$ satisfying Assumptions 6-7 and assuming the validity of the law of large numbers, then it follows from routine analysis that $GIC(M(\theta))=P_{n}\{W(x,p_{M(\theta)})\}$ converges uniformly in probability to its population version $GIC_{\infty}(M(\theta))=P\{W(x,p_{M(\theta)})\}$. It follow from Assumption 7 that $GIC_{\infty}(M_{k}(\theta_{k}))$ is a Lipschitz-continuous, as a function of the parameter $\theta_{k}$, hence it attains its maximum value, denoted by $\mathcal{M}_{k}$, owing to the compact parameter space assumption (aka Assumption 6). Suppose $k_{0}$ is the smallest $k$ such that $M_{k}$ contains the true model. Then, $\mathcal{M}_{k}<\mathcal{M}_{k_{0}}$ for all $1\leq k<k_{0}$, whereas $\mathcal{M}_{k}=\mathcal{M}_{k_{0}}$ otherwise. Note that from Proposition 3 in Cheng & Tong (2024), $\mathcal{M}_{k_{0}}=H_{G}(p^{*})>0$ where $p^{*}$ is the true population pdf. Therefore, if for any $k$, $C(n,k)\to 1$ as $n\to\infty$, then the maximum MIC model selection criterion will not select any $k<k_{0}$, in probability.
Henceforth, consider the case that $k\geq k_{0}$. Since $\mathcal{M}_{k_{0}}>0$, $GIC(k)$ is positive, in probability, i.e., $GIC(k)>0$ holds with probability approaching 1 as sample size increases without bound. For ease of exposition, we shall assume that $GIC(k)$ is positive. Let $k>k_{0}$ be fixed. Consider the increment $D=\log\{MIC(k)\}-\log\{MIC(k_{0})\}=\log C(n,k)-\log C(n,k_{0})+\log GIC(k)-\log GIC(k_{0})$. By Proposition 1, $n\times\{\log GIC(k)-\log GIC(k_{0})\}$ converges weakly to some non-negative distribution. Consequently, $D=\log C(n,k)-\log C(n,k_{0})+O_{p}(1/n)$ so that if $n\times\log\{C(n,k)/C(n,k_{0})\}\to$-$\infty$ as $n\to\infty$, $D$ is negative in probability for $k>k_{0}$. This completes the proof of the consistency of the proposed MIC model selection criterion. ∎